! Required information Exercise 10-11 Effects of Changes in Profits and Assets on Return on Investment (ROI) [LO10-1] [The foRequired: 1. Compute the Springfield clubs return on investment (ROI). (Do not round intermediate calculations. Round your a2. Assume that the manager of the club is able to increase sales by $79,000 and that, as a result, net operating income incre3. Assume that the manager of the club is able to reduce expenses by $3,160 without any change in sales or average operating4. Assume that the manager of the club is able to reduce average operating assets by $50,000 without any change in sales or n
Answer #1
1) ROI = 18960/100000 = 18.96%
2) ROI = (18960+6241)/100000 = 25.20%
3) ROI = (18960+3160)/100000 = 22.12%
4) ROI = 18960/50000 = 37.92%